OpenStudy (anonymous):

Assume a 150 pound person burns 120 calories after 2 hours of sleep and 240 calories after 4 hours of sleep. Part 1 (2 points): Use calories as the y-coordinate and hours as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data Part 3 (2 points): How many calories will this person burn after 8 hours of sleep? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.#

OpenStudy (anonymous):

i really need help on this not sure how to solve it :/

OpenStudy (mathstudent55):

Part 1 You need ordered pairs written as (h, c), where h = hours of sleep and c = calories burned. You are given information for two such oredered paitrs: (2, 120) (4, 240) The slope is difference in y divided by difference in x. In this case, it's difference in c divided by difference in h: slope = (240 - 120)/(4 - 2) = 120/2 = 60 Since you divided c (calories burned) by h (hours of sleep), the significance of the slope is how many calories you burn per hour of sleep.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

just got one more question what would be the answer to part 3?